OpenStudy (rainbow_rocks03):

What is the surface area of the square pyramid? A. 60 cm2 B. 132 cm2 C. 228 cm2 D. 288 cm2

OpenStudy (rainbow_rocks03):

@owllover101

OpenStudy (mathstudent55):

The total surface area is the sum of the area of the base (a square) and the 4 sides (triangles).

OpenStudy (mathstudent55):

For the square, use the formula \(A_{square} = s^2\) For the sides use the formula \(A_{triangle} = \dfrac{bh}{2} \)

OpenStudy (anonymous):

Surface area of any shape is given by Curved Surface area+Area of base. Here, Curved Surface area= (Perimeter of Base*Slant Height)/2=24*8/2=96 Area of base= Side*Side=6*6=36. Total surface area= 96+36=132

OpenStudy (mathstudent55):

Since there are 4 triangular sides, remember to multiply the area of a side by 4 and add that to the area of the square base.

OpenStudy (rainbow_rocks03):

oh

OpenStudy (rainbow_rocks03):

so my answer is 132

OpenStudy (mathstudent55):

\(A_{square} = (6~cm)^2 = 36~cm^2\) \(A_{triangle} = \dfrac{6~cm \times 8~cm}{2} = 24~cm^2\) Area of 4 sides = 24 cm^2 * 4 = 96 cm^2 Total surface area: 36 cm&^2 + 96 cm^2 = 132 cm^2 Yes, you are correct. Don't forget the units, cm^2